### v1

@@ Abstract
Despite the many recent achievements in developing and deploying social robotics, there are still many underexplored environments and applications for which systematic evaluation of such systems by end-users is necessary. While several robotic platforms have been used in gerontological healthcare, the question of whether or not a social interactive robot with multi-modal conversational capabilities will be useful and accepted in real-life facilities is yet to be answered. This paper is an attempt to partially answer this question, via two waves of experiments with patients and companions in a day-care gerontological facility in Paris with a full-sized humanoid robot endowed with social and conversational interaction capabilities. The software architecture, developed during the H2020 SPRING project, together with the experimental protocol, allowed us to evaluate the acceptability (AES) and usability (SUS) with more than 60 end-users. Overall, the users are receptive to this technology, especially when the robot perception and action skills are robust to environmental clutter and flexible to handle a plethora of different interactions.
@@ 1 Introduction
Social robots are not yet commonly found in our public spaces, despite this vision being an imminent reality over 25 years ago [1]. In addition to classic robotic skills, such as object avoidance during navigation, social robots must be able to seamlessly communicate with multiple people through natural verbal and non-verbal interaction. Over the past decade, social robots have been tested in museums, airports, libraries, shopping malls, bars, and hospitals [2, 3, 4, 5, 6, 7], reporting many positive findings. They have been used to successfully make sports and rehabilitation exercises more entertaining [8], assist older adults in care facilities [9, 10] and hospitals [11, 12], navigate and engage with people in public spaces (like concert halls, hotel lobbies, and shopping malls) [9, 13, 5], and engage in multi-party interaction [3, 14]. However, today’s social robots are far from perfect as they are often run in a wizard-of-oz setup (with the researchers controlling the robot’s navigation and dialogue manually) [9, 10, 14]. Those that do function independently are stationary [13, 8], have limited dialogue capacity (rule-based or closed-domain) [3, 5], or are designed for single user interactions that cannot be guaranteed in public spaces.
@@
Such spaces require more complex robotic skills and introduce new underexplored challenges [15]. The robot must be able to fuse multi-sensory input to sense people and objects [16], tracking their positions as the robot navigates through the space [17], all while avoiding collisions. The robot must be able to hear its interlocutors [18] despite background noise, the robot’s ego-noise, the acoustics of the room, and multiple people speaking at the same time. It must understand where people are looking [19], and determine if they are getting frustrated with the robot to aid communication. The robot must move its head and eyes to look at its addressee or nod, move its arms appropriately when pointing, and move in the correct direction when guiding people. Typical speech systems are created to converse with only one individual at a time (e.g. Amazon Alexa, or Siri [20]), whereas pairs and groups of individuals may approach the robot together. People talk to each other as well as to the robot in a multi-party conversation, and the robot’s spoken dialogue system must be able to handle this [21]. Similarly, the navigation skills of the robotic platforms must be able to adapt to new environments without requiring extensive retraining or adaptation.
@@
To tackle the above challenges, the EU’s H2020 SPRING project [^1] https://spring-h2020.eu/ aims to develop social robots that can communicate in complex and unstructured public spaces. The SPRING (Socially Pertinent Robots IN Gerontological healthcare) project is a consortium of five international research labs (computer science & engineering), two industry partners, and importantly, a hospital with research facilities. Our experimental setting is the Broca gerontology day-care hospital, where patients visit when they are suspected to have dementia. Patients typically spend full days at the hospital with a friend or family member for support. The hours are filled with multiple appointments, but a large portion of the day is also spent waiting for test results or the next appointment. Our goal is to provide a system that is both practically useful and also entertaining, to provide participants with information and some light distraction from their otherwise stressful day. The research staff at the hospital run the experiments with volunteer patients, their companions, and the ARI robot (Figure 1).
@@
One of the objectives of the SPRING project, and the main research question of the study presented in this paper, is whether or not patients of a gerontology day-care hospital and their companions will find a social robot useful and acceptable in this environment. Our study is unique in the sense that (i) the evaluation is conducted with real patients and companions of and in a day-care gerontological hospital (ii) with a full-sized humanoid robot (ARI is $1.65$ m tall) (iii) that is capable of advanced sensing and acting skills allowing it to converse with patients and companions. Our findings suggest that both patients and companions find such technology useful and acceptable, especially if it is robust to noise and clutter, and flexible to the plethora of different situations it can encounter.
@@
To be able to conduct this study, we have developed a complex and feature-rich software architecture that is described in Section 2, where we detail the overall system architecture and provide technical details of each robotic ‘skill’, and the improvement beyond the state-of-the-art. Some of the modules in the architecture were not evaluated in the hospital, but only in our laboratories due to Ethical constraints. Where appropriate, we refer to technical reports or publications to condense this paper without it becoming over-simplistic. In Section 3 we describe the experimental setup in the hospital. We discuss the ethical considerations, the protocol, the performance metrics, the main results, and the failure cases. Finally, we conclude and list open topics in Section 4.
@@ 2.1 Overview & Robot Platform
To accomplish its objective of developing a socially pertinent robot, the SPRING consortium developed a novel software architecture for humanoid robots. The architecture is composed of eight modules (Figure 2) responsible for perception (self-localisation, human localisation, speech processing, human behaviour analysis, person manager) and action (experimenter interface, multi-party conversation, non-verbal behaviour generation) processes. They are developed on top of the ROS 1 Noetic [^2] http://wiki.ros.org/noetic middleware. In total, all modules consist of 52 ROS nodes communicating through more than 170 ROS topics and actions. Where relevant, the architecture uses standard ROS messages. In particular, the SPRING architecture is one of the first to fully adopt the REP-155 ‘ROS4HRI’ [^3] https://www.ros.org/reps/rep-0155.html [22] standard for human-robot interaction to combine the different perception modalities (voices, faces, bodies) into a consistent representation of persons that can be used by downstream action modules.
@@
The developed architecture is deployed on the PAL ARI humanoid robot, designed for use as a socially assistive companion [23]. ARI is 1.65m tall, with a differential-drive mobile base, a touch-screen on the torso, movable arms to gesture, and a head with LCD eyes that enable expressive gazing behaviours. It is equipped with a four-microphone array (front torso), an RGB camera (head), and two ${}^{\circ}$ fish-eye cameras (chest and back) allowing us to capture and record the audio and video of the whole interaction from the robot’s perspective. An additional RGB-D camera is located in the front for self-localisation. The robot verbalises given responses using Acapela Text-To-Speech. [^4] https://www.acapela-group.com/ The following sections provide details into the key features of each module.
@@ 2.2 Self-Localisation
The self-localisation module takes the RGB-D and Fisheye images as input and provides a 2D occupancy map indicating obstacles and the current position of ARI in the map. This information is crucial for obstacle avoidance during the navigation of ARI (Section 2.8). The developed module integrates RTABMap [24], a library for mapping and real-time tracking that incorporates ORB-SLAM [25], with an augmented version of the Hierarchical-Localisation (HLOC) [26] algorithm. RTABMap and HLOC provide two distinct, pre-aligned maps that share the same global coordinate system. The pre-alignment is done by aligning the camera centers registered in these maps. The module fuses both information to provide a single consistent map for other modules. This multi-layered approach is crucial for robust and precise localisation in complex and dynamic environments.
@@
Combining state-of-the-art image retrieval and pose estimation techniques, this complex methodology results in state-of-the-art localisation accuracy. Using front and backwards-facing images captured from two consecutive time points (i.e., four fisheye images), the system achieves localisation within the maps of the Broca gerontology day-care hospital with less than 1cm positional error and 0.2 degrees rotational error in 90% of queries. Comparatively, with a single image, the success rate drops to 15% for positional accuracy within 1cm and 30% for rotational accuracy within 0.2 degrees. These findings highlight the enhanced precision of the multiview approach over single-image HLOC localisation used to initialise the real-time camera pose tracking performed by ORB-SLAM.
@@ 2.3 Human Localisation
The main goal of this module is to localise the humans in the space through time. To that aim, we use audio-visual data – more precisely the front fisheye images and the microphone array signals – to detect people, track them over time, and re-identify them (meaning that there is a one-to-one correspondence between persons and tracks). This information is exploited by the Conversation Manager (Section 2.7) to have a time-consistent identifier for each conversation partner and the human-aware navigation to move naturally around humans (Section 2.8). On the audio side, we employed a time-difference of arrival (TDOA) estimation for the audio modality, utilising an instantaneous version of the generalized cross-correlation phase-transform (GCC-PHAT[29]. The GCC-PHAT algorithm reliably provides TDOA readings in frames with a single speaker. This implementation utilises two horizontal microphones from the Respeaker microphone array embedded in ARI. The TDOA readings are then translated into direction of arrival (DOA) estimates using geometric considerations On the visual side, we have implemented a state-of-the-art multi-person visual tracker known as fair multi-object tracking (FairMOT) [30]. This method combines the detection and the re-identification abilities and is based on the well-known residual neural network (ResNet34) [31]. However, the standard architectures are trained for regular cameras, while ARI’s camera has a fisheye lens. This required an annotation and retraining procedure described in [32]. Another important property of FairMOT is that the tracking is based on the Kalman filter, plus an additional matching step that associates detections (from the current time step) and Kalman predictions (from the previous time step), by means of a detection-to-prediction distance matrix. By computing the distance between DOA and the predictions of the Kalman filter, we can seamlessly incorporate audio detections to the tracking pipeline, thus achieving multi-modal speaker detection and tracking.
@@ 2.4 Speech Processing
Regarding the short-term attributes, the activity can be directly obtained from the Sep-TFAnetVAD network, which incorporates an voice activity detector (VAD). The activity patterns of the separated speakers can serve for diarization in the downstream dialogue manager. Short-term identification relies partly on spatial information derived from the current scene. This involves employing a late fusion mechanism that combines visual-based and audio-based DOA estimation, as explained in Section 2.3.
Regarding the identity, we utilise Nvidia’s ECAPA-TDNN model [38] to extract voice-specific (speaker) embeddings, producing a 192-dimensional voice signature vector. The speaker identification module then stores these embeddings in an internal database. The identification process occurs by comparing the cosine similarity between an active speaker and an entry in the database, triggering a match when the similarity exceeds a specified threshold. Finally, the speaker embedding obtained from the speaker extraction algorithm [37] may also serve as a speaker ID.
@@
Transcribing audio is crucial in social robotics, as this component will generate the words comprehended by the robot. We carried out an extended experimental evaluation of several automatic speech recognition (ASR) systems [39] and concluded that the best option was to use Nvidia’s on-prem solution called RIVA (Version 2.7). Apart from demonstrating performance comparable to existing cloud services, RIVA offers the advantage of adaptability to our data distribution. Moreover, it operates on-premises, mitigating potential concerns associated with data and privacy issues.
@@ 2.5 Human Behaviour Analysis
This section describes the modules of ARI regarding the perception of human behaviour, namely: gaze target detection, detection of social acceptance of ARI, and multimodal emotion recognition. These modules input the position and groups of people from Section 2.3, the speech as processed in Section 2.4 as well as the face camera image, and could have direct applications in understanding the people’s intentions, emotional state, and engagement with the robot. While we did not have the opportunity yet to exploit the outcome of these modules in hospital settings, they are part of the software architecture and publicly available as the other ones.
@@
Figure 3: (left) Illustrative example of a relations graph generated by the perception modules. Numerical values are likelihoods that two features belong to the same person. (right) The person manager partitions the graph to maximise overall likelihood. One anonymous person had to be created for the features Body 1 and Voice 2, as they would otherwise be dangling. Note that the Person 1 - Voice 1 edge, for example, does not exist in the original data. It is a computed edge, used to ensure stability if Body 2 is not detected any more and disappears.
@@ 2.7 Multi-party Conversation Manager
Commercial and research spoken dialogue systems (SDSs), conversational agents, and social robots have been designed with a focus on dyadic interactions. That is a two-party conversation between one individual user and a single system/robot. Such interactions can only be guaranteed in specific settings, like people interacting with Siri on their phone, or with Amazon Alexa in single-occupant homes. When Alexa is in a family home, its lack of multi-party capabilities is apparent [60], but this becomes a critical limitation when deploying social robots in public spaces [2, 3, 4, 5, 6, 7], where multi-party conversations (MPCs), involving people talking to both the robot and each other, do commonly occur.

### v2

@@ Abstract
Despite the many recent achievements in developing and deploying social robotics, there are still many underexplored environments and applications for which systematic evaluation of such systems by end-users is necessary. While several robotic platforms have been used in gerontological healthcare, the question of whether or not an autonomous social interactive robot with multi-modal conversational capabilities will be useful and accepted in real-life facilities is yet to be answered. This paper is an attempt to partially answer this question, via two waves of experiments including patients and companions in a day-care gerontological facility in Paris with a full-sized humanoid robot endowed with social and conversational interaction capabilities. The software architecture, developed during the H2020 SPRING project, together with the experimental protocol, allowed us to evaluate the acceptability (AES) and usability (SUS) with more than 60 end-users. Overall, the users are receptive to this technology, especially when the robot’s perception and action skills are robust to environmental clutter and flexible to handle a plethora of different interactions. We also report and discuss some concerns and general comments of the users.
@@ 1 Introduction
Social robots [1, 2], i.e. autonomous machines designed to interact and communicate with humans using social cues like voice, gestures, and facial expressions, are not yet commonly found in our public spaces, despite this vision being an imminent reality over 25 years ago [3]. In addition to classic robotic skills, such as object avoidance during navigation, social robots must be able to seamlessly communicate with multiple people through natural verbal and non-verbal interaction. Over the past decade, social robots have been tested in museums, airports, libraries, shopping malls, bars, and hospitals [4, 5, 6, 7, 8, 9, 10], reporting many positive findings. They have been used to successfully make sports and rehabilitation exercises more entertaining [11], assist older adults in care facilities [12, 13] and hospitals [14, 15], navigate and engage with people in public spaces (like concert halls, hotel lobbies, and shopping malls) [12, 16, 7], and engage in multi-party interaction [5, 17]. However, today’s social robots are far from perfect as they are often run in a wizard-of-oz setup (with the researchers controlling the robot’s navigation and dialogue manually) [12, 13, 17]. Those that do function independently are stationary [16, 11], have limited dialogue capacity (rule-based or closed-domain) [5, 7], or are designed for single user interactions that cannot be guaranteed in public spaces.
@@
Such spaces require more complex robotic skills and introduce new underexplored challenges [18]. The robot must be able to fuse multi-sensory input to sense people and objects [19], tracking their positions as the robot navigates through the space [20], all while avoiding collisions. The robot must be able to hear its interlocutors [21] despite background noise, the robot’s ego-noise, the acoustics of the room, and multiple people speaking at the same time. It must understand where people are looking [22], and determine if they are getting frustrated with the robot to aid communication. The robot must move its head and eyes to look at its addressee or nod, move its arms appropriately when pointing, and move in the correct direction when guiding people. Typical speech systems are created to converse with only one individual at a time (e.g. Amazon Alexa, or Siri [23]), whereas pairs and groups of individuals may approach the robot together. People talk to each other as well as to the robot in a multi-party conversation, and the robot’s spoken dialogue system must be able to handle this [24]. Similarly, the navigation [25] or tracking [26] skills of the robotic platforms must be able to adapt to new environments without requiring extensive retraining or adaptation.
To tackle the above challenges, the EU’s H2020 SPRING project [^1] https://spring-h2020.eu/ aimed to develop social robots that can communicate in complex and unstructured public spaces. The SPRING (Socially Pertinent Robots IN Gerontological healthcare) project is a consortium of five international research labs (computer science & engineering), two industry partners, and importantly, a gerontology hospital with research facilities. In this context, the main research question of the study presented in this paper is whether or not patients of a gerontology day-care hospital and their companions will find an autonomous social robot useful and acceptable in this environment.
@@
More precisely, our experimental setting is the Broca gerontology day-care hospital, where patients visit when they are suspected to have dementia. Patients typically spend full days at the hospital with a friend or family member for support. The hours are filled with multiple appointments, but a large portion of the day is also spent waiting for test results or the next appointment. In order to answer our research question, we have developed a robotic system able to autonomously provide participants with information and some light distraction from their otherwise stressful day, and we have conducted a long series of experiments with volunteer patients and their companions, interacting with the ARI robot, see Figure 1.
@@
The paper has three major contributions. First, a feature-rich robotic software architecture with advanced sensing and acting skills that enable a full-sized humanoid robot (ARI is 1.65 m tall) with autonomous social interaction capabilities. Second, an experimental protocol to evaluate this technology in a day-care gerontology hospital with real patients and companions, including ethical and practical considerations. Third, a series of experimental waves, involving more than 60 participants, conducted at the Broca hospital, their associated results, comparison and, discussion. Our findings suggest that both patients and companions find such technology useful and acceptable, especially if it is robust to noise and clutter, and flexible to the plethora of different situations it can encounter.
@@
The rest of the paper is structured into two main sections, namely Architecture and Experiments, before reaching the Conclusions. In Section 2, we detail the overall system architecture and provide technical details of each robotic ‘skill’, and the improvement beyond the state-of-the-art. Some of the modules in the architecture were not evaluated in the hospital, but only in our laboratories due to Ethical constraints. Where appropriate, we refer to technical reports or publications to condense this paper without it becoming over-simplistic. Readers mostly interested in the conducted experiments and the associated results are strongly encouraged to read the architecture overview in Section 2.1, then jump to Section 3, where we describe the experimental setup in the hospital, including the ethical considerations, the protocol, the performance metrics, the main results, and the failure cases. Finally, we conclude and list open topics in Section 4.
@@ 2.1 Overview & Robot Platform
To understand whether or not patients of a gerontology day-care hospital and their companions will find a social robot useful and acceptable, we developed and implemented a novel software architecture for humanoid robots. Indeed, to our knowledge, the architecture proposed in this paper is the only one developed for enabling multi-modal social interactions with multiple persons in unscripted conversations. The architecture is composed of eight modules (Figure 2) responsible for perception (self-localisation, human localisation, speech processing, human behaviour analysis, person manager) and action (experimenter interface, multi-party conversation, non-verbal behaviour generation) processes. They are developed on top of the ROS 1 Noetic [^2] http://wiki.ros.org/noetic middleware. In total, all modules consist of 52 ROS nodes communicating through more than 170 ROS topics and actions. Where relevant, the architecture uses standard ROS messages. In particular, our architecture is one of the first to fully adopt the REP-155 ‘ROS4HRI’ [^3] https://www.ros.org/reps/rep-0155.html [27] standard for human-robot interaction to combine the different perception modalities (voices, faces, bodies) into a consistent representation of the persons in the scene and that can be used by downstream action modules.
@@
The developed architecture is deployed on the PAL ARI humanoid robot, designed for use as a socially assistive companion [28]. ARI is 1.65m tall, with a differential-drive mobile base, a touch-screen on the torso, movable arms to gesture, and a head with LCD eyes that enable expressive gazing behaviours. It is equipped with a four-microphone array (front torso), an RGB camera (head), and two 180 fish-eye cameras (chest and back) allowing us to capture and record the audio and video of the whole interaction from the robot’s perspective. An additional RGB-D camera is located in the front for self-localisation. The robot verbalises given responses using Acapela Text-To-Speech. [^4] https://www.acapela-group.com/ To our knowledge, the proposed architecture is unique in the sense that it follows several design requirements associated with our experimental goal: it exploits audio and visual data captured with the robotic sensors on board, it represents a social scene involving multiple people via the verbal and non-verbal behavioural cues, it is capable of deciding which social and dialogue actions are needed depending on the representation of the social scene, it has a usable experimenter interface, and finally all the computations are carried out on the premises. In addition, it is the first architecture effectively exploiting the ROS4HRI standard [27].
Generally speaking, our architecture can be used with other robotic platforms, provided that they have similar sensing and acting capabilities (audio-visual perception, self-localisation-oriented cameras, loudspeakers, and a mobile base). While, in practice, some minor adjustments will always be required, our modular structure allows for changes only in the few ROS modules that are in direct contact with the sensors and actuators, while keeping the rest untouched. The following sections provide details of the key features of each module.
@@ 2.2 Self-Localisation
The self-localisation module takes the RGB-D and Fisheye images as input and provides a 2D occupancy map indicating obstacles and the current position of ARI in the map. This information is crucial for obstacle avoidance during the navigation of ARI (Section 2.8). The developed module integrates RTABMap [29], a library for mapping and real-time tracking that incorporates ORB-SLAM [30], with an augmented version of the Hierarchical-Localisation (HLOC) [31] algorithm. RTABMap and HLOC provide two distinct, pre-aligned maps that share the same global coordinate system. The pre-alignment is done by aligning the camera centres registered in these maps. The module fuses both information to provide a single consistent map for other modules, by aligning the camera centers in both maps. This multi-layered approach is crucial for robust and precise localisation in complex and dynamic environments.
@@
Combining state-of-the-art image retrieval and pose estimation techniques, this complex methodology results in state-of-the-art localisation accuracy. Using front and backwards-facing images captured from two consecutive time points (i.e., four fisheye images), the system achieves localisation within the maps of the Broca gerontology day-care hospital with less than 1cm positional error and 0.2 degrees rotational error in 90% of queries. Comparatively, with a single image, the success rate drops to 15% for positional accuracy within 1cm and 30% for rotational accuracy within 0.2 degrees [34]. These findings highlight the enhanced precision of the multiview approach over single-image HLOC localisation used to initialise the real-time camera pose tracking performed by ORB-SLAM.
@@ 2.3 Human Localisation
The main goal of this module is to localise the humans in the space through time. To that aim, we use audio-visual data – more precisely the front fisheye images and the microphone array signals – to detect people, track them over time, and re-identify [^6] Re-identification is a standard task in computer vision, aiming to understand whether two images are of the same or different persons, and has nothing to do with retrieving personal information of an individual from an image. them (meaning that there is a one-to-one correspondence between persons and tracks). This information is exploited by the Conversation Manager (Section 2.7) to have a time-consistent identifier for each conversation partner and the human-aware navigation to move naturally around humans (Section 2.8). On the audio side, we employed a time-difference of arrival (TDOA) estimation for the audio modality, utilising an instantaneous version of the generalized cross-correlation phase-transform (GCC-PHAT[35]. The GCC-PHAT algorithm reliably provides TDOA readings in frames with a single speaker. This implementation utilises two horizontal microphones from the Respeaker microphone array embedded in ARI. The TDOA readings are then translated into direction of arrival (DOA) estimates using geometric considerations. On the visual side, we have implemented a state-of-the-art multi-person visual tracker known as fair multi-object tracking (FairMOT) [36]. This method combines the detection and the re-identification abilities and is based on the well-known residual neural network (ResNet34) [37]. However, the standard architectures are trained for regular cameras, while ARI’s camera has a fisheye lens. This required an annotation and retraining procedure described in [38]. Another important property of FairMOT is that the tracking is based on the Kalman filter, plus an additional matching step that associates detections (from the current time step) and Kalman predictions (from the previous time step), by means of a detection-to-prediction distance matrix. By computing the distance between DOA and the predictions of the Kalman filter, we can seamlessly incorporate audio detections to the tracking pipeline, thus achieving multi-modal speaker detection and tracking [39, 40].
@@ 2.4 Speech Processing
Regarding the short-term attributes, the activity can be directly obtained from the Sep-TFAnetVAD network, which incorporates an voice activity detector (VAD). The activity patterns of the separated speakers can serve for diarisation in the downstream dialogue manager. Short-term identification relies partly on spatial information derived from the current scene. This involves employing a late fusion mechanism that combines visual-based and audio-based DOA estimation, as explained in Section 2.3.
Regarding identity, we utilise Nvidia’s ECAPA-TDNN model [46] to extract voice-specific (speaker) embeddings, producing a 192-dimensional voice signature vector. The speaker identification module then stores these embeddings in an internal database. The identification process occurs by comparing the cosine similarity between an active speaker and an entry in the database, triggering a match when the similarity exceeds a specified threshold. Finally, the speaker embedding obtained from the speaker extraction algorithm [45] may also serve as a speaker ID.
@@
Transcribing audio is crucial in social robotics, as this component will generate the words comprehended by the robot. We carried out an extended experimental evaluation of several automatic speech recognition (ASR) systems in French [47] and concluded that the best option was to use Nvidia’s on-prem solution called RIVA (Version 2.7). Apart from demonstrating performance comparable to existing cloud services, RIVA offers the advantage of adaptability to our data distribution. Moreover, it operates on-premises, mitigating potential concerns associated with data and privacy issues.
@@ 2.5 Human Behaviour Analysis
This section describes the modules of ARI regarding the perception of human behaviour, namely: gaze target detection, detection of social acceptance of ARI, and multimodal emotion recognition. These modules input the position and groups of people from Section 2.3, the speech as processed in Section 2.4 as well as the face camera image, and could have direct applications in understanding the people’s intentions, emotional state, and engagement with the robot. While we have not had the opportunity yet to exploit the outcome of these modules in hospital settings, they are part of the software architecture and publicly available as the other ones.
@@
Figure 3: Example of the processing of social features by the person manager. Faces, voices, bodies, persons, and groups are denoted by different colours. Edge numbers represent the likelihood of two nodes being associated. The relations graph generated by the perception module is shown on the left, and the changed graph by the person manager to maximise the overall likelihood on the right. Gray edges on the left are discarded and purple edges and nodes on the right are added by the person manager. For instance, Person 3 was added for features Body 1 and Voice 2 to be associated with a person. Additional edges were added to ensure stability over time (e.g. robustness against undetected bodies, faces, or voices).
@@ 2.7 Multi-party Conversation Manager
Commercial and research spoken dialogue systems (SDSs), conversational agents, and social robots have been designed with a focus on dyadic interactions. That is a two-party conversation between one individual user and a single system/robot. Such interactions can only be guaranteed in specific settings, like people interacting with Siri on their phone, or with Amazon Alexa in single-occupant homes. When Alexa is in a family home, its lack of multi-party capabilities is apparent [68], but this becomes a critical limitation when deploying social robots in public space [4, 5, 6, 7, 8, 9], where multi-party conversations (MPCs), involving people talking to both the robot and each other, do commonly occur.